# Character is Destiny: Can Large Language Models Simulate Persona-Driven Decisions in Role-Playing?

### v1

@@ 5 Analysis
Raw text Model desc. Human desc.
+Model retr. +Model retr.
GPT-4 - 58.92 70.94
human 92.01 65.82 72.78
Table 4: Results of the human evaluation. The inputs provided to humans are the raw text, model description + model retrievals, and human description + model retrievals, respectively.
Restoration Level Raw input Entity replacement
completely consistent 27.5% 0%
partially consistent 47.5% 37.5%
entirely different 25.0% 62.5%
Table 5: Analysis of Data Leakage Mitigation through Entity Replacement. Three restoration levels represent the model’s ability to generate the original plot before and after replacement, and the percentages represent the proportions of different restoration levels.
In the experiments, we wish to answer three research questions: RQ1) Can LLMs make decisions based on historical data? RQ2) What influences the decision-making of LLMs? RQ3) How can the decision-making of LLMs be improved?
@@ 5.1 Can LLMs make decisions based on historical data?
How do models make decisions based on historical data?
@@
Table 3 reports the accuracy results of baseline methods on the future prediction task in our dataset. Several observations can be made: First, the performance of the combined methods outperforms the use of a single one, suggesting that both holistic and detailed data of characters are important in final decision-making. Second, when provided with a gold explanation, the accuracy consistently exceeds 90%, indicating the rationality of these explanations in the data. Third, the performance gap among different LLMs is not significant while reasoning the answer. This indicates that the main factor for the result is the generated profile, rather than reasoning ability.
How do humans make decisions based on historical data?
We invite three native English-speaking university students to participate in a test where we select six novels they have never heard of before. Each novel has between 5 to 10 characters and their corresponding multiple-choice questions. We test three groups, allowing them to make choices after reading all content before the decision nodes: model-generated description + embedding-retrieved memories or human description + embedding-retrieved memories, as shown in Table 4. The results indicate that humans have a strong ability to make choices after reading the original text. When provided with the compressed content, human performance is slightly better than GPT-4. This indicates a considerable scope for enhancing our method, particularly in acquiring information from the original text that impacts the characters.
Do LLMs know the plot following the node in advance?
Data leakage poses a significant challenge for our benchmarks since the novel data we utilize might have appeared in the pre-training set. We aim for the model to forget relevant plots as much as possible. To this end, we adopt the entity replacement strategy. During the evaluation, we replace character names with placeholders like [character A]. Further, we replace placenames and some distinctive phrases that could lead to data leakage, such as the iconic sentence “For you, a thousand times over” from the novel “The Kite Runner”. We utilize GPT-4 to identify these potential leakages and have applied rules to replace them.
To test the effectiveness of this approach, we input the human summary before and after entity replacement, and ask the model to output the subsequent content, evaluating its capability to restore the plot from the book. The levels of restoration are categorized as “completely consistent”, “partially consistent”, and “entirely different.” We conduct a manual evaluation of 40 characters from 20 books. Specific evaluation details can be found in Appendix B.2. As shown in Table 5, After replacing the entities, the model no longer produces content that is “completely consistent”, and the overall situation of data leakage is significantly alleviated.
@@ 5.2 What influences the decision-making of LLMs?
Figure 3: Heatmap of the impact of motivation types on the results. The results are predicted from the model-generated description, the model-retrieved memory, the combination of both, and CharMap. The reasoning model uses GPT-4.
Does the type of motivation affect the model’s decision-making?
In line with the motivation types presented in Section 3.2, we examine how different types of motivation influence characters’ decision-making. For profiles, we evaluate three scenarios: using only model-generated descriptions, embedding-retrieved memories, and a combination of both. For reasoning, we use GPT-4 uniformly. The results are shown in Figure 3. We find that for all methods, tasks requiring coherent reasoning, such as puzzles and mysteries, are not well answered. This might be because these questions need multi-step reasoning and details from various memories. Moreover, plot-driven questions have lower accuracy when only descriptions are used for the profile. Conversely, character-driven questions are challenging to answer when relying only on memories. We believe this is because character summaries in descriptions better capture the overall essence of the characters, while memories provide direct access to relevant events.
@@
Does the genre of the novel affect the model’s decision-making?
@@
We use the genre tags from novels on the website to analyze the accuracy of character selection across different genres. We conduct experiments on the combination of human description + embedding-retrieved memories, and the reasoning model using GPT-4. As depicted in Figure 4, the accuracy of science fiction, fantasy novels, and romance novels is quite high. This could be due to the characters in these novels often being stylized or having fixed creative patterns and archetypes. In contrast, crime and mystery novels perform poorly, which might be because they involve complex logical chains, and characters in these novels frequently take abnormal actions. Further details about each genre and the complete table can be found in Appendix A.1.
Are expert-annotated motivations considered gold interpretable motivations?
We need to ensure that the problems and motivations are corresponding. As shown in Table 3, when we use these motivations as part of the input, the accuracy of the QA task exceeds 85%. This demonstrates that the motivations effectively aid the model in making selections. In addition, we adopt the evaluation method used in recent research Liu et al. (2023); Li et al. (2024), utilizing GPT-4 to evaluate whether the model-generated profiles contain human motivations. We assign scores from 1 to 3 for the model in this regard:
Score 1: The paragraph fails to convey the core information of the target sentence, exhibiting significant misunderstandings or omissions.
Score 2: The paragraph conveys the main information of the target sentence but misses some details or has minor inaccuracies.
Score 3: The paragraph completely and accurately conveys all the key points of the target sentence. We calculate average scores using different methods.
As Table 3 shows, a higher score suggests that the model-generated profile more accurately represents the information from the manually created motivations. Moreover, higher scores correspond to increased final accuracy, which reiterates the validity of the manual motivations. To uphold the quality of the GPT-4 evaluation, we select a sample of data for human evaluation and comparison and compute the Pearson Correlation Coefficient Cohen et al. (2009) across both evaluation types. The value reaches 0.81, indicating that GPT-4’s evaluation capabilities on this task are on par with humans. The specific evaluation process can be found in Appendix B.3.
@@
Figure 5: Analysis of whether character selection will change. The x-axis represents the input length relative to the node truncation.
Do characters at different times make the same choices?
@@
If faced with the decisions of years past at this moment, would you make the same choices? We conduct a study on this matter. Specifically, we randomly sample 40 characters, half character-driven and half plot-driven. We split the content preceding the decision points into five equal sections and used these various content lengths as input. We conduct experiments on the combination of human description + embedding-retrieved memories, and the reasoning model uses GPT-4. As shown in Figure 5, in the early stages, the accuracy of most characters’ decisions is close to random (25%), potentially due to insufficient information. As more information becomes available, the characters’ decisions tend to be closer to the correct choice. For character-driven decisions, accuracy tends to be stable. For plot-driven, the accuracy rate may change abruptly. This could be due to the relatively stable characteristics of a character, while some sudden events may greatly influence the final choices of the character.
Figure 6: An overview of CharMap, a two-step scenario-specific character profile building approach. It locates content in memory that is more consistent with the current scenario through the connection between description and question.
5.3 How can the decision-making of LLMs be improved?
Based on the conclusion in section 5.2, the model often performs poorly in answering questions requiring multiple details or multi-hop reasoning, possibly due to the overly dispersed memory retrieved. We propose the CHARacter MAPping Profile Synthesis (CharMap), constructing a more scenario-specific profile in two steps. As shown in Figure 6, first, after obtaining the description generated by the model, we input it along with the question into the model, asking the model to locate the plot in the description relevant to the current scene based on the question. Second, we use these episodes as queries to retrieve related memory, and then input them into the inference model together with the description. This can take advantage of the overall description of the character storyline in the description, thereby better retrieving related memories. As shown in Table 3, the accuracy after using CharMap is 6.01% higher than directly concatenating the description with memory. As seen from Figure 3, the accuracy rate of each category of questions improves after using the CharMap, especially the Puzzles category that requires multi-hop reasoning.
@@ 6 Conclusion
In this work, we propose NextDecisionPrediction, the first task to evaluate the persona-driven decision-making of LLMs, testing whether LLMs can reconstruct actual storylines through historical character data. We construct a dataset LifeChoice, including 1,401 characters from 396 books and their significant decisions. We analyze the performance of LLMs on this task from various aspects, such as novel genres, character motivations, and model solution approaches. Based on this, we propose a method, CharMap, which achieves better results by constructing a character profile more relevant to the current scene in two steps.
Decision-making is an important and complex behavior of humans, and we hope to explore the possibility of LLMs making personal decisions. This could establish important evaluation criteria for role-playing models and Personal LLM assistants.
@@
The partial evaluation method we proposed is dependent on GPT-4, which could be biased towards GPT-4 generations. Furthermore, although we have alleviated the data leakage problem through entity replacement, there is still a small amount of data that may have data leakage, especially in some classic novels. Finally, our dataset is constructed through the decision of high-quality novel characters. However, compared to human choice, this part of the data is not sparse or challenging enough. We hope to construct real human decision-making data while ensuring privacy.
Appendix A Dateset Details
A.1 Categories of novel
@@
Magical Realism Novels: Magical realism stories take place in the real world but have characters who take magical elements for granted. These magical elements do not exist in real life, but they are perfectly normal in the realm of magical realism.
A.2 Categories of motivations
@@
Appendix B Manual Annotation
@@
For all individuals involved in the annotation, we provide compensation based on the local minimum hourly wage.
B.1 Manual Filtering Rules
This is a supplement to Section 3.1. After constructing the multiple-choice question data using GPT-4, we perform manual filtering.
For each annotator, we provide novel summaries and character analyses written by human literature experts on the Supersummary website. Each annotator is asked to score the questions constructed by GPT-4 based on the following evaluation criteria:
1. Comprehensiveness
Rule 1.1: Evaluators must ensure that each multiple-choice question fully considers the character’s background, context, and motivation. The questions should reflect the true decisions and experiences of the character within the narrative.
Scoring Guide:
2 points: The question is detailed and comprehensive, aligning perfectly with the character’s background and motivation.
1 point: The question aligns generally but is missing key aspects of the character’s background information or motivational nuances.
0 points: The question significantly misaligns with the character’s background or motivation.
2. Logical Consistency
Rule 2.1: Evaluators should assess the internal consistency and plausibility of the question within the narrative thread. The content and structure of the multiple-choice question must be consistent with the plot and the character’s logical decision-making process.
Scoring Guide:
2 points: The question is entirely consistent with the character’s known decisions and the structure of the plot.
1 point: The question is generally consistent but has minor inconsistencies in detail.
0 points: The question is logically inconsistent with the character’s known decisions or the structure of the plot.
3. Challenge Level
Rule 3.1: Evaluators need to assess the plausibility of the incorrect options. Wrong options should be reasonably believable and attractive within the constraints of the character’s background and motivations, making the questions sufficiently challenging.
Scoring Guide:
2 points: All incorrect options are highly plausible, convincingly misleading.
1 point: Most incorrect options are reasonable, but one or two lack plausibility.
0 points: Incorrect options are obviously illogical and lack the ability to mislead.
4. Alignment with Character Motivation
Rule 4.1: Evaluators must assess whether the question correctly guides the testing model to step into the role and make a choice, i.e., testing if the model can replicate the real storyline’s choices. It is crucial that the character’s motivations, as articulated by literary experts, are a central component reflected in these questions.
Scoring Guide:
2 points: The question unambiguously points to a specific character decision point, accurately testing the model’s ability to role-play.
1 point: The question points to a character decision point to some extent, but the indicators are not clear enough, potentially reducing the accuracy of the model’s role-playing test.
0 points: The question fails to clearly define the character decision point, unable to effectively test the model’s role-playing ability.
Additional Notes:
1. Before starting the evaluation, each evaluator must understand the core motives and development axes of the character by reading summaries and analyses of the novels created by literary experts.
2. Ensure that evaluators are familiar with all background material before scoring any questions.
3. Evaluators should reference the analyses by literary experts of the characters to evaluate each of GPT-4’s multiple-choice questions, maintaining consistency of standards.
4. Application of the evaluation rules should be flexible and adapted to the specific context; scoring standards may be adjusted for special cases.
We evaluated the scores of each annotator and only retained the data with an average score of more than 6 points.
B.2 Manual Evaluation of Data Leak
Below are the manual evaluation standards for data leakage issues. We invited three annotators to evaluate the role stories continued by the model, with three levels:
Completely Consistent: The output is deemed "completely consistent" if the model restores the plot with high accuracy, which includes correct character interactions, significant plot points, and events as they happen in the book, despite the replacement of character names and other distinctive elements.
Partially Consistent: The output is "partially consistent" if the model produces content that partially matches the original plot, meaning it can still restore some elements of the story correctly, but there are noticeable discrepancies in the sequence of events, character behaviors, or outcomes not aligning with the text as it is in the book.
Entirely Different: The output is considered "entirely different" if the model’s content bears no resemblance to the original plot, character interactions, or events from the book after the entity replacement. This output suggests that the model is generating new content without relying on potential memorized data from its pre-training set.
The annotators need to select the appropriate level of story restoration for each character based on these three situations. The final proportion of each restoration level is shown in Table 5.
B.3 Manual Evaluation of GPT-4 Scores
In this section, we use GPT-4 to score whether the character profile generated by the model contains the manually annotated motivation. To ensure the quality of GPT-4 assessment, we sample 40 pieces of data and invite 3 native English-speaking college students for manual evaluation. We provide the annotators with the same scoring standards as GPT-4, and each annotator scores all data. We calculate the Pearson Correlation Coefficient Cohen et al. (2009) between the model and manual scores. The final results show that the Pearson Correlation Coefficient between GPT-4 and each annotator exceeds 0.7, and the correlation coefficient with the average score of annotators is 0.81. This indicates that GPT-4’s evaluation ability on this task can be comparable to that of humans.

### v2

@@ 5 Analysis
Raw text Concat. CharMap
GPT-4 - 63.05 68.10
human 92.01 66.82 74.78
Table 4: Results of the human evaluation. Concat. refers to the direct concatenation of Description and Memory.
In the experiments, we wish to answer two research questions: RQ1) Can LLMs make decisions based on historical data? RQ2) What influences the decision-making of LLMs?
@@ 5.1 Can LLMs Make Decisions Based on Historical Data?
Analysis of Model Results
@@
Table 3 presents the accuracy results of different RPLA methods on the LifeChoice. Additionally, we evaluate the results when the model is provided with gold motivation, and several observations can be made: First, the method that uses both Description and Memory surpasses the one that uses only one, suggesting that both holistic and detailed data of key characters are essential in final decision-making. Second, when gold motivation is provided, the accuracy consistently exceeds 80%, indicating the rationality of these motivations in the data. Third, the performance gap among different LLMs is not significant while reasoning the answer. This indicates that the main factor for the result is the generated profile rather than reasoning ability. Last, CharMap outperforms the method that directly concatenates Description and Memory by 5.03%, proving its effectiveness. This scenario-specific profile better assists RPLA in decision-making.
Humans are Good Decision-makers
We invite three native English-speaking university students to take a test in which we select six novels they have never heard of before. Each novel has between 3 to 5 characters and their corresponding multiple-choice questions. We provide each person with three data sets for each key character in two books: the full original text before the decision point, direct concatenation Description and Memory result, and the result from CharMap. As shown in Table 4, compared to direct concatenation, the CharMap results are easier for humans to understand. Additionally, humans slightly outperform GPT-4 in reasoning answers based on the profiles, indicating that humans can understand subtle character decisions better than models. When given the raw text, humans can achieve an accuracy rate of 92.01%, suggesting there is still significant room for improvement in RPLA methods.
Figure 4: The impact of the number of book reviews on accuracy in LifeChoice, with new books being those not present in the training corpus of LLMs.
Mitigation and Analysis of Data Leakage
Data leakage is a significant challenge since our data might appear in the model’s pre-training corpus. During the data collection phases in section 3.1, we adopt various preventive measures. For evaluation, we employ an entity replacement strategy, substituting character names, locations, and other entities with placeholders. We believe data leakage relates to the amount of relevant corpus used during LLM pre-training, with more popular books having more related corpus. To verify this, we use the number of reviews on the book review website [^3] https://www.douban.com/ to indicate a book’s popularity and evaluate the results of books with different review counts on LifeChoice. We use CharMap to build profiles and GPT-4 as the role-playing model, sampling thirty books with different numbers of reviews, including thirty books not in the LLMs’ corpus (published after November 6 for gpt-4-1106-preview). As shown in Figure 4, the model’s accuracy significantly improves when the number of reviews exceeds 5,000. In contrast, books with fewer than 5,000 reviews show slight fluctuation and results similar to those not in the LLMs’ corpus. Therefore, it can be considered that for books with a low number of reviews, data leakage has little impact on CharMap. In section 3.1, we use 5,000 reviews as a threshold to filter the books.
Analysis of Long-Context LLMs
Long context is an essential feature of LifeChoice, and directly using long-context models for role-playing is an exciting topic. Making decisions based on extensive context tests a model’s ability to understand global data and reason from a character’s perspective. We evaluate two long-context models: Claude3-sonnect and kimi-chat. As shown in Table LABEL:tab:longcontext, although the performance of long-context models is not as strong as CharMap, they still demonstrate potential in role-playing. LifeChoice, as a task requiring multiple reasoning points and an overall understanding of the context, can also serve as a vital benchmark for evaluating long-context models.
@@ 5.2 What Influences the Decision-making of LLMs?
Figure 5: Heatmap of the impact of motivation types on the results. The results are predicted from the Incremental updating, the embedding-retrieved memory, the direct concatenation of both, and CharMap. The role-playing model uses GPT-4.
The Impact of Motivation Types
In line with the motivation types presented in Section 3.2, we examine how different types of motivation influence characters’ decision-making. For profiles, we evaluate four methods: the Incremental updating, the embedding-retrieved memory, the direct concatenation of both, and CharMap. For reasoning, we use GPT-4 uniformly. The results are shown in Figure 5. We find that tasks requiring coherent reasoning, such as puzzles and mysteries, are not well answered for all methods. This might be because these questions need multi-step reasoning and details from various memories. Moreover, plot-driven questions have lower accuracy when descriptions are used only for the profile. Conversely, character-driven questions are challenging to answer when relying only on memories. We believe this is because character summaries in descriptions better capture the overall essence of the characters, while memories provide direct access to relevant events.
@@
The Impact of Novel Genres
@@
We use the genre tags from novels on the website to analyze the accuracy of character selection across different genres. We conduct experiments on the the direct concatenation of description and memories, and the role-playing model using GPT-4. As depicted in Figure 6, the accuracy of science fiction, fantasy novels, and romance novels is quite high. This could be because the characters in these novels are often stylized or have fixed creative patterns and archetypes. In contrast, crime and mystery novels perform poorly, which might be because they involve complex logical chains, and characters in these novels frequently take abnormal actions. Further details about each genre and the complete table can be found in Appendix B.1.
@@
Figure 7: Analysis of whether character selection will change. The x-axis represents the input length relative to the point truncation.
The Impact of Temporal Data
@@
If faced with the decisions of years past at this moment, would you make the same choices? We conduct a study on this matter. Specifically, we randomly sample 40 characters, half character-driven, and half plot-driven. We split the content preceding the decision points into five equal sections and used these various content lengths as input. We conduct experiments on the combination of human description + embedding-retrieved memories, and the role-playing model is GPT-4. As shown in Figure 7, in the early stages, the accuracy of most characters’ decisions is close to random (25%), potentially due to insufficient information. As more information becomes available, the characters’ decisions tend to be closer to the correct choice. For character-driven decisions, accuracy tends to be stable. For plot-driven, the accuracy rate may change abruptly. This could be due to the relatively stable characteristics of a character, while some sudden events may greatly influence the final choices of the character.
@@ 6 Conclusion
In this work, we propose the first task to evaluate the decision-making of RPLAs, testing whether LLMs can accurately reconstruct decisions using historical data. We construct LifeChoice, which includes 1,462 characters from 388 books and their life choices. Extensive experiments on LifeChoice demonstrate the promising performance of RPLAs in decision simulation. Additionally, we propose CharMap, which uses persona-based memory retrieval to enhance decision-making. We hope this work provides better evaluation benchmarks for RPLAs and directs the future development of personal LLM assistants.
@@
In this paper, we primarily investigate whether fictional characters can recreate their choices within a book. Although we have controlled the quality of the novels, there may still be issues with the plot and characters since the author designed the storyline, which can result in illogical choices within the book. Furthermore, as our research focuses mainly on fictional characters, there is a certain gap compared to real-world humans. For example, the author fictionalizes some story backgrounds, which may impact the model’s generation results.
Ethics Statement
Use of Human Annotations
Our institution recruits annotators to implement the annotations of motivation recognition dataset construction. We ensure the privacy rights of the annotators are respected during the annotation process. The annotators receive compensation exceeding the local minimum wage and have consented to use motivation recognition data they process for research purposes. Appendix C provides further details on the annotations.
Risks
The LifeChoice dataset is sourced from novels and analysis data written by human literary experts. However, we cannot guarantee that this data is entirely free from toxic or discriminatory language. Additionally, our data is generated by GPT-4, which may introduce some inherent biases and hallucinations of the model.
Appendix A Prompts
In this section, we provide the key prompts we used, including prompts for selecting decision points, locating the node’s position, constructing multiple-choice questions, system prompts for role-playing characters, and the prompt of CharMap.
A.1 Selecting Decision Points
As mentioned in section 3.1, the first step in constructing our data is selecting the character’s Decision Points. In this step, our input data consists of all raw data from Supersummary, including the current character’s description, all chapter summaries, and book analysis. We provide this data to GPT-4, requesting it to output three types of data: the character’s decision point, the corresponding gold motivation, and the chapters related to this choice. Table 5 presents the character’s description, Table 6 presents the chapter summaries, and Table 7 presents the book analysis, these examples are all from the 2024 novel "A Calamity of Souls.". Table 8 illustrates the prompt for selecting decision points.
A.2 Locating the Node
Furthermore, for the character’s decision points, we provide the previously identified decision points and their corresponding chapters from the original book to GPT-4. This allows it to precisely determine the position in the original book that should be segmented, helping to avoid data leakage. Table 9 shows the prompt for locating.
A.3 Constructing Multiple-Choice Questions
After selecting the Decision Points, our next step is constructing Multiple-Choice Questions. In this step, our input data consists of all the input and output data from the previous step(Appendix A.1). We ask GPT-4 to construct a multiple-choice question regarding the character’s decision based on this data, outputting the scenario in which the character is situated, the question, and four options. Table 10 shows the prompt for constructing multiple-choice questions.
A.4 System Prompts for Role-Playing
The prompt for role-playing as the character can be found in Table 11.
A.5 Prompts for CharMap
The prompt for CharMap can be found in Table 12.
Appendix B Dateset Details
B.1 Categories of novel
@@
Magical Realism Novels: Magical realism stories take place in the real world but have characters who take magical elements for granted. These mystical elements do not exist in real life, but they are perfectly normal in magical realism.
B.2 Categories of motivations
@@
Appendix C Manual Annotation
This is a supplement to Section 3.1. After constructing the multiple-choice question data using GPT-4, we perform a manual examination. For each annotator, we provide key character descriptions, chapter summaries, and book analyses written by human literature experts on the Supersummary. Each annotator is asked to score the questions constructed by GPT-4 based on the rules shown in Table 13. We evaluated the scores of each annotator and only retained the data with an average score of more than 6 points.
@@ Appendix C Manual Annotation
We provide compensation based on the local minimum hourly wage for all individuals involved in the annotation.
Appendix D Future direction
Building personal agents for everyone is an exciting topic. We have explored how fictional characters can determine their subsequent actions based on historical data, proposing possibilities for combining role-playing with personalized models. We believe there are additional directions to explore:
  • Real-life version of LifeChoice The behavior of fictional characters often stems from the author’s design, which can lead to logical inconsistencies. In contrast, real-world human behavior data do not have this issue. How to construct real human historical data and related behavior data is a question worth exploring.
  • Improving RPLA performance in life choices Although CharMap has achieved decent results, better methods are needed to balance reasoning efficiency (which depends on the length of input tokens) while achieving superior outcomes.
  • More complex downstream decision tasks The decisions we select are often significant choices for fictional characters, resulting in a large decision space without a fixed task framework. Identifying more systematic tasks by integrating social sciences is a challenge that needs to be addressed in the future.